OpenStudy (anonymous):

The sum of two consecutive odd integers is 124, What are the integers?

OpenStudy (anonymous):

@wwhitlock

OpenStudy (johnweldon1993):

Let our first integer be 2n + 1 Why 2n + 1? because any number times 2 comes out even right? and then adding 1 makes it automatically odd.. so we have 2n + 1......and the next consecutive number would be 2n + 3 since it would just be the same number plus 2 we know that equals 124....so \[\large (2n + 1) + (2n + 3) = 124\] \[\large 4n + 4 = 124\] \[\large 4n = 120\] \[\large n = 30\] if 'n' = 30...we have \[\large 2(30) + 1 = 61\] \[\large 2(30) + 3 = 63\] These are our 2 numbers

OpenStudy (wolf1728):

2x+1 + 2x+3 = 124 4x=120 x=30 2x1+1 = 61 the other number is 63

OpenStudy (wwhitlock):

wolframalpha.com agrees. Medal should go to @johnweldon1993

OpenStudy (wolf1728):

okay - he was quicker (and more detailed)

OpenStudy (johnweldon1993):

It shouldn't because I gave out the answer -_- gotta watch that! lol...but this is the method you would use to solve these types of problems

OpenStudy (wolf1728):

Sheesh 2 medals ? Well I'll send you one too.